Let A and B be events such that P (A) = 0.7, P (B) = 0.6 and P (AUB) = 0.8. Find P (A∩B) and P (A’∩B).  


1
Expert's answer
2021-11-23T17:59:17-0500
"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"P(A\cap B)=P(A)+P(B)-P(A\cup B)"

i)

"P(A\cap B)=0.7+0.6-0.8=0.5"

ii)


"P(A'\cap B)=P(B)-P(A\cap B)=0.6-0.5=0.1"